OpenStudy (anonymous):

Write an expression for the apparent nth term of the sequence. (Assume that n begins with 1.) -3-(4/1),-3-(4/2),-3-(4/3),-3-(4/4),-3-(4/5)

OpenStudy (anonymous):

an = -3 -4/n correct me if I'm wrong ^_^

OpenStudy (anonymous):

Well That is one of my answer choices... is it possible you could explain how you got it so that I can do the other problem like it please?

OpenStudy (anonymous):

or help me with my other one?

OpenStudy (anonymous):

sure, first write it down on paper and tell me what number do you see is changing?

OpenStudy (anonymous):

you'll notice that -3 and -4 stay the same and that only the denominator of 4 is increasing by one

OpenStudy (anonymous):

the denominator

OpenStudy (anonymous):

exactly and you said assume n starts from n =1

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so, you can conclude that the formula for this sequence is as follows: an = -3 -4/n for n to be any number :)

OpenStudy (anonymous):

clear now? ^_^

OpenStudy (anonymous):

oh ok so n is the number that's changing yes makes sense thanks

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

if I work out the problem like it could you check my work?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Write the nth term of the arithmetic sequence as a function of n. \[a _{1}=4, a _{k+1}=a _{k}-5\] I think the answer is: \[a _{n}=4-5n\]

OpenStudy (anonymous):

Write the nth term of the arithmetic sequence as a function of n. \[a _{1}=8, a _{k+1}=a _{k}+7\] I think the answer is: \[a _{n}=8+7n\]

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

for both of them!?!

OpenStudy (anonymous):

I guess, yes :)

OpenStudy (anonymous):

Sweet Thank you so much for helping me!

OpenStudy (anonymous):

you are most welcome ^_^! good luck!

OpenStudy (anonymous):

but I'm not really sure of the answers though, they look good :)

OpenStudy (anonymous):

lol well Ijust turned in the test and I got both those wrong:( lol still got a B but oh well lol

OpenStudy (anonymous):

hmmm, since ak = 8, and ak+1 = 7, the value is decreasing so I think it's: an = 8 - n

OpenStudy (anonymous):

oops >_< my bad! I'm sorry you got them wrong :(

OpenStudy (anonymous):

so if n = 1, then an = 8 -1 = 7

OpenStudy (anonymous):

I'm really sorry, I should've used my head >_< at first. Dearest apologies!

OpenStudy (anonymous):

That's ok no biggy the last one I posted the answer is 1+7n

OpenStudy (anonymous):

Ha ha don't sweat it it's all good

OpenStudy (anonymous):

close? ^^"

OpenStudy (anonymous):

alright, next time I'll make sure to be accurate! :)

OpenStudy (anonymous):

ha ha sounds like a plan stan either way you helped me thanks!

OpenStudy (anonymous):

Have a good rest of your day!

OpenStudy (anonymous):

np ^_^

OpenStudy (anonymous):

you too :)

OpenStudy (anonymous):

and angoo is still quite , LOL! are you sick or something AG?

OpenStudy (anonymous):

I am not sick, I am very busy thinking on how to make some money online

OpenStudy (anonymous):

You'll lose your hair at a very young age .

OpenStudy (anonymous):

lol Good luck with that! STart your own tutoring service lol

OpenStudy (anonymous):

ttt..

OpenStudy (anonymous):

who cares

OpenStudy (anonymous):

well, YOU should care

OpenStudy (anonymous):

I've noticed that the more you think the more moody you get lol, ease up a lil =P

OpenStudy (anonymous):

in the end, money doesn't make a person happy :)

OpenStudy (anonymous):

you get, you spend, finito~

OpenStudy (anonymous):

you are right

OpenStudy (anonymous):

it is not money I care about, it is friends

OpenStudy (anonymous):

exactly :)

OpenStudy (anonymous):

if I were you, I'll go visit some friends now and hang out for a bit ^_^

OpenStudy (anonymous):

my goal now it to find a VERY good friend

OpenStudy (anonymous):

it'll be EXTRAVAGANT! LOL

OpenStudy (anonymous):

friends are everywhere, the one you choose depends on how much you like that person ^_^

OpenStudy (anonymous):

How do you find friends with money? Materialism will only get you so far and for the most part greedy friends

OpenStudy (anonymous):

i dont have many friends, most ppl in my area are busy drinking and clubbing, I dont drink and dont go to clubs so..

OpenStudy (anonymous):

i dont make money to get friends

OpenStudy (anonymous):

Friends, are those who seek your comfort and support. Those people are hard to find but soon to get :)

OpenStudy (anonymous):

don't lose hope =P

OpenStudy (anonymous):

right on! That's awesome Same here!

OpenStudy (anonymous):

don't stress about it , and you'll see how simple ppl will come into your life ^_^

OpenStudy (anonymous):

Good luck! YOu'll find friends just be sure not to over look someone who might be your friend:)

OpenStudy (anonymous):

simply*

OpenStudy (anonymous):

that's right mmb :)

OpenStudy (anonymous):

well, ppl come easily, but not all of those ppl have right intentiond, some dont even care about the real person in me

OpenStudy (anonymous):

then those are not your friends

OpenStudy (anonymous):

I have a simple question for you, "Do you want Friends?" answer with a "yes/no" only

OpenStudy (anonymous):

i have girls asking me out all the time (yes it it sounds crazy, but for some reason they do it), but I feel they just care about the looks and not the real me

OpenStudy (anonymous):

YES, GOOD friends

OpenStudy (anonymous):

not fake short term ones

OpenStudy (anonymous):

That's it! you've stated what you have wanted, then you're done, all you've got to do is be patient and they'll come + stop thinking abt it now :) and move on ^_^ with your life

OpenStudy (anonymous):

those are called acquaintances... There will be a lt of those I have a lot of them anyways but have you ever thought that spome of those acquaintances come to you for a certain reason, like there is something that attracts them to you... you're nice, funny, they like being around you part of it is leading by example, listening eventually they catch on and you both earn respect for eachother

OpenStudy (anonymous):

you said what you've wanted, and you'll get it, stop thinking abt it and move on ahead. Simple as that ~

OpenStudy (anonymous):

LOL, acquaintances, LOL LOL LOL

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

ok, lets leave the job to the law of attraction :)

OpenStudy (anonymous):

exactly! :)

OpenStudy (anonymous):

Wallace was smart

OpenStudy (anonymous):

do the happy dance now and you'll feel better lol, kidding, but really you'll feel better ^_^

OpenStudy (anonymous):

ha ha that's what I call them and refer to them as that too lol works for me

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

That's not a joke sstarica! I just did a happy dance about the score on my math test lol it makes you feel better and trust me my morning has been rough! lol

OpenStudy (anonymous):

I always do the happy dance and trip in the end, I always do and still don't know WHY! .___.

OpenStudy (anonymous):

>_<

OpenStudy (anonymous):

Ha ha dance off's happy dances basically just the greatest thing ever invented

OpenStudy (anonymous):

yep ~ :)

OpenStudy (anonymous):

but tripping is not ^^"

OpenStudy (anonymous):

well, believe it or not but I got some music on...

OpenStudy (anonymous):

and i am about to dance a dance

OpenStudy (anonymous):

YAY! let's kill the night with a dance ~ :D

OpenStudy (anonymous):

I enjoy dancing ~

OpenStudy (anonymous):

LOL :D

OpenStudy (anonymous):

It's pretty subzero right!

OpenStudy (anonymous):

you got me excited! >_<

OpenStudy (anonymous):

I am a horrible dancer but man if there is something to groove too.... lol

OpenStudy (anonymous):

....I feeeeel good....

OpenStudy (anonymous):

Whoo Hooo!!!

OpenStudy (anonymous):

ooooooh yeeeaahhh~

OpenStudy (anonymous):

mimi, you are smiling so wide I can see it from here

OpenStudy (anonymous):

=D lol drop by anytime mmb and I'll teach you how to dance, though it's not something you learn, just comes from the heart :)